You have just received a windfall from an investment you made in a friends business. She will be paying you $12,230 at the end of this year, $24,460 at the end of next year, and $36,690 at the end of the year after that (three years from today). The interest rate is 3.6% per year. a. What is the present value of your windfall? b. What is the future value of your windfall in three years (on the date of the last payment)?
Answer #1

a.Present value of inflows=cash inflow*Present value of discounting factor(rate%,time period)

=12230/1.036+24460/1.036^2+36690/1.036^3

=$67591.18(Approx).

b.We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=12230*(1.036)^2+24460*(1.036)+36690

=$75156.97(Approx).